OpenStudy (anonymous):

Molar Volume of a Gas lab: How do I determine the volume of hydrogen collected at STP and SATP, and the molar volume at STP and SATP?

OpenStudy (anonymous):

Room temperature (°C) 22.0 Barometric pressure (kPa) 100.5 Water vapor pressure at 22 °C (kPa) 2.60 Mass of Mg ribbon (g) 0.0500 Volume of hydrogen gas (mL) 48.3 Observations of reaction • when the Mg came into contact with the HCl bubbles of gas (H2) were produced • the H2 produced displaced the water in the graduated cylinder

OpenStudy (anonymous):

oh here: 2HCl(g) + Mg(s) -> H2(g) + MgCl2(aq) nMg = 0.0020571899 mol nH2 = 0.0020571899 mol Wet H2 pressure = 100.5 kPa Dry H2 pressure = 97.9 kPa --- Vstp = 22.4 L/mol, and Vsatp = 24.8 L/mol Vstp:0 degrees Celsius and 101.325 kPa while Vsatp: 25 degrees Celsius and 100 kPa. 100.5 kPa x 760 mm Hg/101.325 kP = 753.8119911 mmHg So to determine the volume of hydrogen gas collected at STP: V2 = (753.8119911 mmHg)(48.3 mL)(273 K) / (295 K)(760 mmHg) = 44.33403003 mL Molar volume at STP: mol/L = 44.33403003 mL / 0.0020571899 mol = 21550.77176 mL/mol = 21.55077176 L/mol To determine volume of hydrogen gas collected at SATP: V2 = (100.5 kPa)(48.3 mL)(298 K) / (295K)(100kPa) = 49.03514237 mL Molar volume at SATP: mol/L = 49.03514237 mL / 0.0020571899 mol = 23835.98246 mL/mol = 23.8359824 L/mol Are my answers correct? :$

OpenStudy (anonymous):

You have the work out with: Volume of H2 (STP) : V' = ( 48.3 ml * 97 .9 * 273 ) / 100.5 * 295 = 43.54 mL Mole of H2 = .0434 ml / .0021 = 21.17 mol

OpenStudy (anonymous):

so wait only the volume of H2 (ST) is wrong?

OpenStudy (anonymous):

The key is you need to use corrected Pressure 97 .9 kPa, because we want the dry H2 gas!

OpenStudy (anonymous):

ohh didn't I use it?

OpenStudy (anonymous):

ohhh.. i don't use the wet one?

OpenStudy (anonymous):

V2 = (753.8 mmHg)(48.3 mL)(273 K) / (295 K)(760 mmHg) = 44.33 mL I doesn't look like you use the dry one !!!

OpenStudy (anonymous):

so basically I replace the 100.5 kPa from both of the calculations STP and SATP to 97.9 kPa?

OpenStudy (anonymous):

lol, thank you!

OpenStudy (anonymous):

ohh it's 20993.23937.. o.O

OpenStudy (anonymous):

The 20.73 mol is the correct one.

OpenStudy (anonymous):

wait I am so confused now D:

OpenStudy (anonymous):

this part: Mole of H2 = .0435 L / .0021 = 20.73 mol

OpenStudy (anonymous):

ohhh i see it. Never mind, thanks again!

OpenStudy (anonymous):

Volume of H2 (STP) : V' = ( 48.3 ml * 97 .9 * 273 ) / 101.3 * 295 = 43.197 mL Molar Volume of H2 = .043197 ml / .0021 = 20.99 mol